4x(x-1)-5x(x)=3 solve for x

please help and thankyou

Answer by nyc_function(2741)

I will begin and you can finish.
4x(x - 1)- 5x(x)= 3
4x^2 - 4x - 5x^2 = 3
-x^2 - 4x = 3
-x^2 - 4x - 3 = 0
Can you do it now?
http://nycfunction.webs.com/